OpenStudy (anonymous):

I'm having a little bit of trouble with logs (Well, more specifically, rewriting logs in terms of others). Here is one problem: log 3 (10) = 2.1 log 3 (11) = 2.2 log 3 (8) = 1.9 Without using a calculator, find log 3 (330)

OpenStudy (anonymous):

\[\log_{3} 330=\log_{3} \left( 3*11*10 \right)=\log_{3} 3+\log_{3} 11+\log_{3} 10=\frac{ \log 3 }{\log 3 }+2.2+2.1\]=? \[(\frac{ \log x }{ \log x }=1)\]

OpenStudy (anonymous):

Ah, okay. That's what I was thinking. The log 3 (8) sorta threw me off. :)

OpenStudy (doc.brown):

I hate the term log, I always say\[log_2(8)=?\]two to the power of what equals 8 or in your case \[log_3(330)=?\]three to the power of what is 330?